OpenStudy (anonymous):

The sum of 3 consecutive odd internees is -15. What are the numbers? A. {-3, -1, 1} B. {-7, -9, -11} C. {-4, -5, -6} D. {-3, -5, -7}

OpenStudy (anonymous):

Intergers***

Directrix (directrix):

Intergers*** should be Integers. There is only one "r" in Integer. Let x represent the first of the three consecutive odd integers. x + 2 is the second and x + 4 is the third. Solve this equation for x: x + x + 2 + x + 4 = -15 Post what you get and we can compare answers, okay?

OpenStudy (anonymous):

What number would be x ?

Directrix (directrix):

To know that, you need to solve this equation: x + x + 2 + x + 4 = -15 x + x + 2 + x + 4 = -15 3x + 6 = -15 For the next step, subtract 6 from both sides.